would you please explain to me how to solve the problem, 4m - n = m, for ,m ?
-thanks

Answer by stanbon(75887)

would you please explain to me how to solve the problem,
4m - n = m, for ,m ?
----------------
Subtract "m" from both sides to get:
3m -n = 0
Add "n" to both sides to get:
3m = n
Divide both sides by 3 to get:
m = n/3
=============
Cheers,
Stan H.